OpenStudy (anonymous):

find the vector with the given polar coordinates :

OpenStudy (anonymous):

\[a) r= 4 , \theta=\frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

i cant approach the answer :'(

OpenStudy (anonymous):

Polar to rectangular coordinate transform: \((x, y) = (r \cos \theta, r \sin \theta)\)

OpenStudy (anonymous):

okay i can do this part , how do i get the vector

OpenStudy (anonymous):

I'm assuming you need a vector in \(\mathbb{R}^2\). Vectors in \(\mathbb{R}^2\) are just ordered pairs of real numbers, in other words you can express the vector in terms of its real number components.

OpenStudy (anonymous):

So the vector is simply \((r \cos \theta, r \sin \theta) = (4 \cos \frac{\pi}{3}, 4 \sin \frac{\pi}{3})\)

OpenStudy (anonymous):

really? that is so simple , i have been absence alot from my calculus classes

OpenStudy (anonymous):

that will be a relief

OpenStudy (anonymous):

it cant be assumed in R^3 right?

OpenStudy (anonymous):

@Alchemista